The average hour spent in computer by 20 senior high school students during online class is 7 hours with a standard deviation of 2 hours. Construct a 98% confidence interval of the average hour spent by all senior high school students. Identify the width of the interval.


1
Expert's answer
2022-03-21T06:35:24-0400

"\mu=7 \plusmn z_{alpha/2}\frac{\sigma}{\sqrt{n}}=7 \plusmn 2.33 \frac{2}{\sqrt{20}}=7 \plusmn 1.04"

width of interval is 1.04x2=2.08